OpenStudy (anonymous):

Help please! 63w^15x^15 divided by 7wx^7

jhonyy9 (jhonyy9):

so 63 divided by 7 do you know how many is ,yes ?

jhonyy9 (jhonyy9):

w^15 / w = w^(15-1) = ? and x^15 / x^7 = x^(15-7) = ? hope this will help you